You construct a portfolio out of 4 assets. The betas of the assets are 1, 0.5, 1.9, and 0.9, and the respective weights of the assets in your portfolio are 20%, 15%, 35%, and 30%.

What is the beta of your portfolio? (please round to 2 decimal places )

If the expected market return the next year is 9% and the risk-free rate is 4%, what is the expected return of your portfolio that year? (please round to the nearest percent )

%
Answer #1

1)

Beta = 0.2*1 + 0.15*0.5 + 0.35*1.9 + 0.3*0.9

Beta = 0.2 + 0.075 + 0.665 + 0.27

Beta = 1.21

2)

Expected return = Risk free rate + beta(market return - irks free rate)

Expected return = 0.04 + 1.21(0.09 - 0.04)

Expected return = 0.04 + 0.0605

Expected return = 0.1005 or 10.05%